write the dimension of A/B in the given relation p=(A-T)^2/Bx, where p=pressure,x=distance, t=time?

Grade:11

1 Answers

Aamina
7 years ago
The dimensions of A are M0L0T1  , because the quantities of same dimensions can be added or subtracted .
 
P = (A-T)2/ Bx      [ given equation ]
 
dimensions of pressure = [ML-1 T-2 ]
 
dimensions of time and distance are [T] and [L] respectively.
 
By substituting the dimensions in the given equation,
 
M1L-1T-2 = T2 / BL
 
[B] =  T2/M1L-1T-2L
 
[B] = M-1L0T4
 
Dimensions of A/B =  M0L0T1/M-1L0T4  = M1L0T-3 = MT-3  .